OpenStudy (anonymous):

t^2-t/6=35/6 solve the equation check your solution show work

OpenStudy (anonymous):

t^2-t/6-35/6=0 factor 1/6 (2 t-5) (3 t+7) t = -7/3 t = 5/2

OpenStudy (anonymous):

www.calc-tech.com